Β Evaluate (π‘«πŸ + πŸ”π‘« + πŸ—)π’š = 𝒆 βˆ’πŸ‘π’• βˆ’ πŸ‘π’„π’π’”πŸ’π’• + π’π’π’ˆ5


1
Expert's answer
2021-10-04T16:47:59-0400

Β Let us solve the differential equation "(D^2 + 6D + 9)y = e^{−3t} − 3\cos 4t +\log 5."

The characteristic equation "k^2+6k+9=0" is equivalent to "(k+3)^2=0," and hence has the roots "k_1=k_2=-3." It follows that the general solution is of the form "y=(C_1+C_2t)e^{-3t}+y_p,"

where "y_p=at^2e^{-3t}+b\cos 4t+c\sin 4t+d." Then "y'_p=2ate^{-3t}-3at^2e^{-3t}-4b\sin 4t+4c\cos 4t,
y''_p=2ae^{-3t}-6ate^{-3t}-6ate^{-3t}+9at^2e^{-3t}-16b\cos 4t-16c\sin 4t."

We get the equation "2ae^{-3t}-6ate^{-3t}-6ate^{-3t}+9at^2e^{-3t}-16b\cos 4t-16c\sin 4t+6(2ate^{-3t}-3at^2e^{-3t}-4b\sin 4t+4c\cos 4t)+9(at^2e^{-3t}+b\cos 4t+c\sin 4t+d)=e^{-3t}-3\cos 4t+\log 5,"

which is equivalent to

"2ae^{-3t}+(-7b-24c)\cos 4t+(-7c-24b)\sin 4t+9d=e^{-3t}-3\cos 4t+\log 5."

We conlude that "2a=1,\ -7b-24c=-3,\ -7c-24b=0,\ 9d=\log 5." It follows that "a=\frac{1}2,\ d=\frac{1}9\log 5, c=-\frac{24}7b,\ 7b-\frac{576}7b=3." Consequently, "a=\frac{1}2,\ d=\frac{1}9\log 5, \ b=-\frac{21}{527}, c=\frac{72}{527}."


We conclude that the general solution of the differential equation "(D^2 + 6D + 9)y = e^{−3t} − 3\cos 4t +\log 5" is

"y=(C_1+C_2t)e^{-3t}+\frac{1}2t^2e^{-3t}-\frac{21}{527}\cos 4t+\frac{72}{527}\sin 4t+\frac{1}9\log 5."